Hugh buys $8000 worth of stock in an electronics company which he hopes to sell afterward at a profit. The company is developing a new laptop computer and a new desktop computer. If it releases both computers before its competitor, the value of Hugh's stock will jump to $22,000. If it releases one of the computers before its competitor, the value of Hugh's stock will jump to $16,000. If it fails to release either computer before its competitor, Hugh's stock will be worth only $5000. Hugh believes that there is a 70% chance that the company will release the laptop before its competitor and a 60% chance that the company will release the desktop before its competitor. Find Hugh's expected profit. Assume that the development of the laptop and the development of the desktop are independent events.

A) $12,200 B) $9200 C) $11,300.00 D) $17,200 E) $15,920
Answer #1

P(both before) = 0.7*0.6 = 0.42 and profit = 22000 - 8000 = 14000

P(neither before) = (1-0.7)(1-0.6) = 0.12 and profit = 5000 - 8000 = -3000

P(one of the item before) = 1 - 0.42 - 0.12 = 0.46 and profit = 16000 - 8000 = 8000

E(x) = 14000*0.42 + (-3000)*0.12 + 8000 * 0.46 = $9200

Hence, expected profit is Option B) $9200